what happens to the circumference of a circle if you double the radius? and what happens if you double the diameter? and what happens if you triple the radius?
Found 2 solutions by mananth, stanbon:
Answer by mananth(16949)

The circumference has the formula 2pir OR pi d
where r is the radius or d the diameter.
The circumference is in direct variation with the radius and the diameter
so if r Or d is doubled the circumference will double.
same with tripling

Answer by stanbon(75887)

what happens to the circumference of a circle if you double the radius?
Circum = 2(pi)r
----
If you double r you get 2(pi)(2r) = 2[2(pi)r] ; you double the Circumference
------------------------------------------------
and what happens if you double the diameter?
Circum (pi)d = (pi)(2r)
--If you double d you get (pi)(2d) = 2(pi)d
So, if you double d, you double the Circumference.
------------------------------------------------
and what happens if you triple the radius?
You get: C = 2(pi)(3r) = 3[2(pi)r]
So, if you triple r, you triple the circumference.
-----
Cheers,
Stan H.
=====================